### v1

@@ 6 Summary
We have applied improved methodologies and clean stellar template libraries (Knabel et al. 2025) to obtain new stellar velocity dispersions using spectra obtain at the Keck and VLT Telescopes for the lens galaxies in the SL2S sample (Gavazzi et al. 2012; Ruff et al. 2011; Sonnenfeld et al. 2013a, b, 2015, Sheu et al. 2025, submitted). The SL2S lens sample is very well matched in terms of deflector redshift and Einstein radii to quasar lenses used by the TDCOSMO collaboration for time delay cosmography (Millon et al. 2020), and it can be used as an external dataset (Birrer et al. 2020) to improve constraints on the Hubble constant and other cosmological parameters (Birrer & Treu 2021), provided stellar velocity dispersions are accurately measured. We use these new measurements to investigate the lens mass fundamental plane (Bolton et al. 2007, 2008a; Auger et al. 2010) as a tool to compare the internal structure of massive galaxies. Our main results can be summarized as follows:
  • The new measurements are on average consistent with the old measurements, but there is significant scatter on an individual system basis.
  • Systematic errors are dominated by the choice of stellar templates and are of the order of 2-3%, depending on the SNR of the spectra.
  • Stellar templates introduce positive covariance between velocity dispersion measurements across the galaxy samples. The covariance is at the level of 1-2%, depending on spectral quality, for galaxies observed with the same instrumental setup. It is significantly smaller (sub-percent) for spectra obtained with different instrumental setups.
  • We derive a simple formula relating the slope of the mass density profile to the observed stellar velocity dispersion, Einstein radius, and effective radius under the assumption of a constant virial coefficient, and show that it is consistent within the errors with the observations.
  • The SL2S, SLACS, and TDCOSMO galaxies follow a tight lens mass fundamental plane, consistent within the errors with the one derived for the SLACS galaxies alone.
  • We introduce a new version of the lens mass fundamental plane, relaxing the assumption of isothermal mass density profiles. We find that this new version is also well defined, albeit with larger error bars as expected, and the three samples are mutually consistent within $1\sigma$ uncertainty.
In terms of big picture, two major conclusions of this study are the following. First, with the new kinematic measurements and systematic error estimates, the SL2S galaxies with the best spectra can be used as external datasets for time delay cosmography. Second, the internal mass structure of massive elliptical galaxies evolves with redshift. However, this evolution happens within the lensing mass plane, which sets tight constraints on their evolutionary mechanism, such as dry mergers (Nipoti et al. 2009) or envelope accretion (Nipoti 2025).
@@ Acknowledgements.
The first two authors should be regarded as joint first authors. We are grateful to William Sheu and our friends and colleagues of the TDCOSMO collaboration for many useful discussions on these topics. TT acknowledges support by NSF through grants NSF-AST-1906976 and NSF-AST-1836016, and from the Moore Foundation through grant 8548. AJS received support from NASA through the STScI grants HST-GO-16773 and JWST-GO-2974.
Table 1. For each lens we list the instruments used, SNR per Å of the spectra, measured instrumental resolution, BIC weighted and equally weighted velocity dispersions, and associated statistical and systematic uncertainties. The BIC determined weights used for each galaxy are reported in Table 2. The systematic uncertainties are Bessel corrected. The last column shows the relative error associated with uncertainties in the measurement of instrumental resolution. For galaxies with two spectra of sufficient spectral resolution, we report both values. The one with the highest SNR is listed first, and it is our fiducial measurement.
@@
Table 2: We report the BIC-weighted average statistical uncertainty, the average systematic uncertainty (both Bessel corrected and uncorrected), the average amplitude of the off-diagonal terms of the covariance matrix between elements of the sample (both Bessel corrected and uncorrected) for the whole SL2S sample and subsamples grouped by instrument and SNR. The BIC determined weights given to the two libraries for each subsample, and the number of galaxies in each subsample is also provided.
Sample $<\delta\bar{\sigma}/\bar{\sigma}>$ $<\Delta\bar{\sigma}/\bar{\sigma}>$ $\sqrt{<C_{i,j}\bar{\sigma}/\bar{\sigma_{i}}\bar{\sigma_{j}}>_{i\neq j}}$ $<\Delta_{B}\bar{\sigma}/\bar{\sigma}>$ $\sqrt{<C_{B,i,j}\bar{\sigma}/\bar{\sigma_{i}}\bar{\sigma_{j}}>_{i\neq j}}$ Indo-US MILES Num. of Obj.
SL2S S/N ¿ 20 3.47% 0.66% 0.22% 2.14% 0.72% 0.05 0.95 15
SL2S S/N (13, 20] 4.60% 1.07% 0.65% 1.70% 1.04% 0.27 0.73 18
SL2S S/N $\leq$ 13 7.40% 1.68% 0.90% 3.07% 1.65% 0.81 0.19 14
SL2S All 5.05% 1.38% 0.67% 2.22% 1.08% 0.26 0.74 47
SL2S-LRIS S/N ¿ 20 3.73% 1.39% 1.07% 2.38% 1.83% 0.78 0.22 8
SL2S-LRIS S/N (13, 20] 4.68% 1.42% 0.97% 2.09% 1.42% 0.37 0.63 13
SL2S-LRIS S/N $\leq$ 13 6.90% 0.91% 0.43% 2.73% 1.30% 0.94 0.06 10
SL2S-LRIS All 5.13% 1.63% 1.02% 2.36% 1.48% 0.62 0.38 31
SL2S-Other S/N ¿ 20 3.16% 0.003% 0.001% 1.97% 0.45% 0.001 0.999 7
SL2S-Other S/N (13, 20] 4.38% 0.40% 0.04% 0.71% 0.06% 0.19 0.81 5
SL2S-Other S/N $\leq$ 13 8.68 % 2.49% 1.55% 3.69% 2.30% 0.35 0.65 4
SL2S-Other All 4.94% 0.10% 0.02% 1.96% 0.35% 0.001 0.999 16
Table 3: Values used in scaling relation calculations and figures. For each dataset (S2LS, SLACS/KCWI, TDCOSMO), we report object name, deflector and source redshift, effective radius, Einstein radius, stellar velocity dispersion corrected to an equivalent aperture of radius equal to half the effective radius, and slope of the mass density profile as inferred from lensing. References are: SL2S (Sonnenfeld et al. 2013a, Sheu et al. 2025; submitted); SLACS(Auger et al. 2010; Tan et al. 2024; Knabel et al. 2024); TDCOSMO (Birrer et al. 2020, and TDCOSMO collaboration 2025, in prep.)
@@
[…]

### v2

@@ 6 Summary
We have applied improved methods and clean stellar template libraries (Knabel_Mozumdar_2025) to obtain new stellar velocity dispersions using spectra that were obtained at the Keck and VLT Telescopes for the lens galaxies in the SL2S sample (Gavazzi12; Ruff11; Sonnenfeld_2013a; Sonnenfeld13b; Sonnenfeld15; William_2025_dinos_2). The SL2S lens sample matches the quasar lenses used by the TDCOSMO collaboration for time-delay cosmography very well in terms of deflector redshift and Einstein radii (Millon_2020_TDCOSMO_I), and it can be used as an external dataset (birrer20) to improve the constraints on the Hubble constant and other cosmological parameters (B+T21), provided stellar velocity dispersions are accurately measured. We used these new measurements to investigate the lensing-mass fundamental plane (Bolton07; Bolton08; Auger_2010) as a tool to compare the internal structure of massive galaxies. Our main results are summarized below.
  • The new measurements are consistent on average with the old measurements, but there is significant scatter for individual systems.
  • Systematic errors are dominated by the choice of stellar templates and are about 2-3%, depending on the S/N of the spectra.
  • Stellar templates introduce a positive covariance between velocity dispersion measurements in the galaxy samples. The covariance is at the level of 1-2%, depending on spectral quality, for galaxies that were observed with the same instrumental setup. It is significantly smaller (sub-percent) for spectra that were obtained with different instrumental setups.
  • We derived a simple formula (Eqs. 17 and 18) that relates the slope of the mass-density profile to the observed stellar velocity dispersion, Einstein radius, and effective radius under the assumption of a constant virial coefficient, and we showed that it is consistent within the errors with the observations.
  • The SL2S, SLACS, and TDCOSMO galaxies follow a tight fundamental plane of the lens mass that is consistent within the errors with the same plane derived for the SLACS galaxies alone.
  • We introduced a new version of the lensing-mass fundamental plane by relaxing the assumption of isothermal mass-density profiles. This new version is also well defined, but the error bars are larger, as expected, and the three samples are mutually consistent within $1\sigma$ uncertainty.
In terms of the larger picture, the two main conclusions of this study are that first, with the new kinematic measurements and systematic error estimates, the SL2S galaxies with the best spectra can be used as external datasets for time-delay cosmography (as discussed by Milestone25). Second, the internal mass structure of massive elliptical galaxies evolves with redshift. This evolution occurs within the lensing-mass fundamental plane, however, which sets tight constraints on their evolutionary mechanism, such as dry mergers (NTB09) or envelope accretion (Nipoti2025).
@@ Acknowledgements.
The first two authors should be regarded as joint first authors. We are grateful to William Sheu and our friends and colleagues of the TDCOSMO collaboration for many useful discussions on these topics. PM, SK, and TT acknowledge support by NSF through grants NSF-AST-1906976, NSF-AST-1836016, and NSF-AST-2407277, and from the Moore Foundation through grant 8548. AJS received support from NASA through the STScI grants HST-GO-16773 and JWST-GO-2974.
Appendix A Measured quantities for SL2S and data from external samples
Table LABEL:table:veldis_measurements. Measured values for the SL2S sample.
@@ Appendix A Measured quantities for SL2S and data from external samples
Notes. For each lens we list the instruments used, S/N per Å of the spectra, measured instrumental resolution, BIC weighted and equally weighted velocity dispersions, and associated statistical and systematic uncertainties. The BIC determined weights used for each galaxy are reported in Table 1. The systematic uncertainties are Bessel corrected. The last column shows the relative error associated with uncertainties in the measurement of instrumental resolution. For galaxies with two spectra of sufficient spectral resolution, we report both values. The one with the highest S/N is listed first, and it is our fiducial measurement.
Table 3: Values used in scaling relation calculations and figures. [^6] For each dataset (S2LS, SLACS / KCWI, TDCOSMO), we report object name, deflector and source redshift, effective radius, Einstein radius, stellar velocity dispersion corrected to an equivalent aperture of radius equal to half the effective radius, and slope of the mass density profile as inferred from lensing. References are: SL2S (William_2025_dinos_2; Sonnenfeld_2013a); SLACS(Auger_2010; T24; Knabel24); TDCOSMO (Milestone25; birrer20)
@@
$\left(\arcsec\right)$ $\left(\arcsec\right)$ $\mathrm{\left(km\ s^{-1}\right)}$